OpenStudy (anonymous):

y = 5x – 4 y = 5x – 5 a. no solutions b. one solution c. infinitely many solutions

OpenStudy (anonymous):

is it b. 1 solution???

OpenStudy (anonymous):

put it in a matrix

OpenStudy (anonymous):

?

OpenStudy (anonymous):

5x - 4 = 5x - 5 5x - 5x = -5 + 4 0x = -1 No solution

OpenStudy (anonymous):

its true.

OpenStudy (anonymous):

no solution y cannot = 5x - more than 1 constant